# Well-clipped cones behave themselves under all finite quotients, the cone conjecture under most

### v1

@@ Abstract.
We use this framework to prove the movable cone conjecture for finite quotients of various projective varieties of Calabi–Yau type, notably products of projective primitive symplectic varieties, abelian varieties, and smooth rational surfaces underlying klt Calabi–Yau pairs. This entails Enriques manifolds. We deduce that such finite quotients admit finitely many unmarked small $\mathbb{Q}$-factorial modifications, and that the nef cone conjecture holds for them.
@@ 1. Introduction
A pair $(X,\Delta)$ is the data of a normal $\mathbb{Q}$-factorial complex projective variety $X$ and of an effective $\mathbb{Q}$-divisor $\Delta$ on $X$ with coefficients in $[0,1]$. A Calabi–Yau pair is a pair $(X,\Delta)$ such that the $\mathbb{Q}$-Cartier divisor $K_{X}+\Delta$ is numerically trivial. A variety $X$ is called of Calabi–Yau type if there exists a klt Calabi–Yau pair $(X,\Delta)$. (For the definition of a klt pair, we refer the reader to [21, Definition 2.34].)
@@
Let $(X,\Delta)$ be a klt Calabi–Yau pair that decomposes as
@@
Our proof starts with an observation: In most instances of pairs for which the cone conjecture has been proved, much more structure was in fact discovered on the cone under study than one may expect: Movable cones of abelian varieties are self-dual homogeneous cones by [38], and so are movable cones of their smooth finite quotients by [31]; Nef cones of K3 surfaces are cut out by the orthogonal hyperplanes to $(-2)$-curve classes in the hyperbolic cone given by the self-intersection form by [43]; Movable cones of projective hyperkähler manifolds are described very precisely by Markman in [29], and so are movable cones of projective primitive symplectic varieties with terminal $\mathbb{Q}$-factorial singularities by [23].
@@
With this terminology, Theorem 1.4 follows from the following result.
@@
Let $(X,\Delta)$ be a pair. Suppose that the movable cone $\mathrm{Mov}(X)$ is well clipped in a self-dual homogeneous cone preserved by the action of $\mathrm{PsAut}(X,\Delta)$, and that the pair $(X,\Delta)$ satisfies the movable cone conjecture. Then the movable cone conjecture holds for every quotient pair $(X/G,\Delta_{G})$ of $(X,\Delta)$ by a finite subgroup $G$ of $\mathrm{Aut}(X,\Delta)$.
@@
Let $\mathscr{C}$ be a well-clipped cone in a self-dual homogeneous cone $\mathscr{A}$ in $V=V_{\mathbb{Z}}\otimes\mathbb{R}$. Let $\Gamma$ denote the group $\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$. Let $G$ be a finite subgroup of $\Gamma$. Then, the invariant cone $\mathscr{C}^{G}$ is well-clipped. Furthermore, if there exists a rational polyhedral fundamental domain for the action of $\Gamma$ on $\mathscr{C}^{+}$, then there exists a rational polyhedral fundamental domain for the action of the centralizer $C_{\Gamma}(G)$ on ${\mathscr{C}^{G}}^{+}$.
@@
Acknowledgements. I want to thank A. Höring, H.-Y. Lin, G. Pacienza and L. Wang for discussions in Fall 2022 that convinced me that descending the Morrison–Kawamata cone conjecture under finite quotients was both non-trivial and interesting. Although they do not appear in the final paper, examples and arguments from B. Skauli’s Master thesis [41] were quite formative to my understanding of this question. I am also thankful to M. Monti, K. Oguiso, A. Quedo and A. Sarti for their interest in this problem and for their questions on earlier communications, and in particular to G. Pacienza for discussions in Mai 2023 in Nancy, to C. Lehn for his helpful comments on a first draft, to S. Filipazzi and N. Tsakanikas for useful remarks on the symplectic setting during a visit to Lausanne in March 2025. I am partially supported by the project GAG ANR-24-CE40-3526-01.
@@ 2. Preliminaries
If $G$ is a finite subgroup of $\mathrm{GL}(V)$, we denote by
@@
the invariant subspace. For a cone $\mathscr{C}$ in $V$ that is preservec by $G$, we denote by $\mathscr{C}^{G}:=\mathscr{C}\cap V^{G}$ the invariant cone.
@@ 2.A. Self-dual homogeneous cones
Let $n\geq 3$. Let $q$ be the quadratic form of signature $(1,n-1)$ on $\mathbb{R}^{n}$, and pick $h\in\mathbb{R}^{n}$ such that $q(h)>0$. The non-degenerate open cone
@@
A classification of $\mathbb{R}$-indecomposable self-dual homogeneous cones was achieved by the 1934 result of Jordan–von Neumann–Wigner [17] on the classification of formally real Jordan algebras, and by the 1970ies Koecher–Vinberg theorem stating an equivalence of categories between formally real Jordan algebras and self-dual homogeneous cones (see e.g. [20], [10, Theorem III.2.1 and Section III.3], and Theorem 2.12 for a precise statement). Here we present the classification.
@@ A Jordan algebra $(V,e,\circ)$ is called formally real if for all $x_{1},\ldots,x_{k}\in V$, we have
The next result is due to Koecher and Vinberg, see for instance [20, Theorem 15].
@@ 2.D. Useful facts from birational geometry
This subsection gathers a few facts from birational geometry, which we will mostly use in the proof of Theorem 1.4. Our first lemma is very close to a result of Xu for the cone $\mathrm{Mov}^{\rm e}$ [47, Theorem 24, Remark on Page 26] and Gachet-Lin-Stenger-Wang for the cone $\mathrm{Nef}^{\rm e}$ [11, Proposition 1.6]. We state it and prove it herefor the cones $\mathrm{Mov}^{+}$ and $\mathrm{Nef}^{+}$, under somewhat weaker assumptions than in [47, Theorem 24].
@@
Note that $F:=f^{*}\mathrm{Mov}^{+}(Y)$ is a face of the cone $\mathrm{Mov}^{+}(X)$. By [11, Proposition 3.6], since $(X,\Delta)$ satisfies the movable cone conjecture, the subgroup $\mathrm{Stab}(F)$ of $\mathrm{PsAut}^{*}(X,\Delta)$ that stabilizes the face $F$ acts on $F$ with a rational polyhedral fundamental domain. To conclude, it suffices to check that $\mathrm{Stab}(F)$ is a subgroup of $f^{-1}\circ\mathrm{PsAut}^{*}(Y,\Delta_{Y})\circ f$.
For $g\in\mathrm{Stab}(F)$, note that there exists $\alpha:X\dashrightarrow X^{\prime}$ a small $\mathbb{Q}$-factorial modification such that the relative interiors of $g^{*}f^{*}\mathrm{Nef}(Y)$ and $f^{*}\alpha^{*}\mathrm{Nef}(Y)$ intersect by [4, Theorem 1.2] (see also [11, Proposition 4.10]). By [11, Lemma 4.2], this means that $f\circ g=\alpha\circ f$, thus $g\in f^{-1}\circ\mathrm{PsAut}^{*}(Y,\Delta_{Y})\circ f$ as wished. ∎
@@
Apply [11, Theorem 1.5 (i), (3) $\Rightarrow$ (4)] to the pair $(X,\Delta)$, and restrict to the Mori chambers in $N^{1}(X)$ associated to small $\mathbb{Q}$-factorial modifications of $Y$ (pre-composed with $f$). ∎
@@
The next lemma is inspired by [9, Lemma 4.6], and is therein also credited to C. Casagrande.
@@
so by [11, Lemma 4.2] (see also [19, Lemma 1.5] in dimension 3), the birational map
@@
Note that applying [9, Lemma 4.6] to the identity from $X$ to itself, any other decomposition of $X$ as a product differs by a mere permutation. Thus, by [15, Exercise III.12.6] and by Lemma 2.17, for a pseudoautomorphism $\alpha$ of $X$, there is a permutation $\sigma$ of the integers between $1$ and $r$ and, for every index $i$, a small $\mathbb{Q}$-factorial modification $\alpha_{i}:X_{i}\dashrightarrow X_{\sigma(i)}$, such that
@@
We conclude with a lemma inspired by [30], [23, Section 7, Page 31], and [11, Proof of Theorem 6.1 (iii) (3a), Page 26-27]. Here, one may think of the $(V_{w})_{w\in W}$ as wall hyperplanes corresponding to flops.
@@
Fix $i$. We now prove the nef cone conjecture for $(X_{i},{\alpha_{i}}_{*}\Delta)$. Denote by $\Sigma_{i}$ the sum of the cones of the form ${\beta_{j}^{-1}}^{*}\Pi\cap\mathrm{Nef}(X_{i})$ for $1\leq j\leq s_{i}$. Note that $\Sigma_{i}$ is a itself rational polyhedral: Indeed, by assumption on the wall hyperplanes $(V_{w})$, any intersection of the form $\alpha_{k}^{*}\mathrm{Nef}(X_{k})\cap\Pi$ for $1\leq k\leq r$ is a rational polyhedral cone. If the translates of $\Sigma_{i}$ by $\mathrm{Aut}(X_{i},{\alpha_{i}}_{*}\Delta)$ cover the cone $\mathrm{Amp}(X_{i})$, then [25, Proposition 4.1] concludes. Let $D\in\mathrm{Amp}(X_{i})$. There exists $\beta\in\mathrm{PsAut}(X,\Delta)$ such that $\beta^{*}\alpha_{i}^{*}D\in\Pi$, in particular it belongs to a chamber $\beta_{j}^{*}\alpha_{i}^{*}\mathrm{Amp}(X_{i})$ for some $j$. By [11, Lemma 4.2], this forces
@@ 3.A. Definition and examples of well-clipped cones
In general, Assumption (ii) could be rephrased, in a lattice-theoretic context, by saying that the $(e_{i})_{i\in I}$ are roots of the lattice $V_{\mathbb{Z}}$, see [5, Section 0.1]. However, we prefer to avoid this terminology altogether, as it is often expected of roots to have norm $-2$, whilst our $(e_{i})_{i\in I}$ can have any negative square.
@@
    The nef cone of a smooth del Pezzo surface $\Sigma$ is rational polyhedral, and well clipped. For $\rho(\Sigma)=1$ or $2$, it follows from Item (1). Otherwise, we can take the quadratic form to be the intersection form on $N^{1}(\Sigma)$, and the $(e_{i})_{i\in I}$ in Remark 3.4 to be the finitely many classes of $(-1)$-curves.
@@
    The movable cone of a smooth projective hyperkähler manifold is well clipped by [28, Lemma 6.22] and [29, Theorem 1.1].
@@
    The movable cone of a projective primitive symplectic variety with terminal $\mathbb{Q}$-factorial singularities is well clipped [23, Theorem 3.10, Lemma 4.6]; [18, Definition 3.1, Theorem 3.11].
@@
    The movable cone of an abelian variety is self-dual homogeneous, thus well clipped [38, Theorem 4.3 and the two paragraphs thereafter].
@@
    The nef cone of $\mathbb{P}^{2}$ blown up at $9$ general points and the movable cone of $\mathbb{P}^{3}$ blown up at 8 very general points are both well clipped, by [33] and [42] respectively.
@@
    with $h^{1}(\mathscr{O}_{X_{i}})=0$ for $2\leq i\leq k$ is well clipped if and only if the movable cone of each factor is well clipped. This follows from [15, Exercise III.12.6], Lemmas 2.16 and 3.5.
@@
By contradiction, assume that $E^{2}\leq-3$. Then $-K_{S}\cdot E>0$. Since $E$ spans an extremal ray of the Mori cone of $S$, the cone theorem gives an elementary Mori contraction $\varepsilon:S\to S^{\prime}$ with exceptional divisor $E$. Computing discrepancies, one checks that $E$ is contracted to a terminal, hence smooth point of $S^{\prime}$. Castelnuovo’s contraction theorem applies and $E$ has square $-1$ in $S$, contradiction. ∎
@@
    By the Beauville–Bogomolov decomposition [3], either $\chi(\mathscr{O}_{\Sigma})\geq 0$ or $S=\Sigma$ is a bielliptic surface. Lemma 3.7 and the fact that bielliptic surfaces have Picard rank $2$ respectively conclude.

### v2

@@ Abstract.
We use this framework to prove the movable Morrison–Kawamata cone conjecture for finite quotients of various projective varieties of Calabi–Yau type, notably products of complex projective primitive symplectic varieties, abelian varieties, and smooth rational surfaces underlying klt Calabi–Yau pairs. This entails Enriques manifolds in the sense of Oguiso–Schröer. We also provide Galois descent statements implying the movable Morrison–Kawamata cone conjecture for abelian varieties over arbitrary perfect fields.
@@ 1. Introduction
A pair $(X,\Delta)$ is the data of a normal $\mathbb{Q}$-factorial complex projective variety $X$ and of an effective $\mathbb{Q}$-divisor $\Delta$ on $X$ with coefficients in $[0,1]$. A Calabi–Yau pair is a pair $(X,\Delta)$ such that the $\mathbb{Q}$-Cartier divisor $K_{X}+\Delta$ is numerically trivial. A variety $X$ is called of Calabi–Yau type if there exists a klt Calabi–Yau pair $(X,\Delta)$. (For the definition of a klt pair, we refer the reader to [22, Definition 2.34].)
@@
Let $(X,\Delta)$ be a klt Calabi–Yau pair defined over $\mathbb{C}$ that decomposes as
@@
Our proof starts with an observation: In most instances of pairs for which the cone conjecture has been proved, much more structure was in fact discovered on the cone under study than one may expect: Movable cones of abelian varieties are self-dual homogeneous cones by [41], and so are movable cones of their smooth finite quotients by [33]; Nef cones of K3 surfaces are cut out by the orthogonal hyperplanes to (2)(-2)-curve classes in the hyperbolic cone given by the self-intersection form by [46]; Movable cones of projective hyperkähler manifolds are described very precisely by Markman in [31], and so are movable cones of projective primitive symplectic varieties with terminal $\mathbb{Q}$-factorial singularities by [25].
@@
With this terminology, Theorem 1.4 follows from the following result applied over the field $k=\overline{k}=\mathbb{C}$.
@@
Let $(X,\Delta)$ be a pair defined over a perfect field $k$. Let $(\overline{X},\overline{\Delta})$ denote their base change to the algebraic closure $\overline{k}$ of $k$. Suppose that the movable cone $\mathrm{Mov}(\overline{X})$ is well clipped in a self-dual homogeneous cone preserved by the action of $\mathrm{PsAut}(\overline{X},\overline{\Delta})$, and that the pair $(\overline{X},\overline{\Delta})$ satisfies the movable cone conjecture. Then the movable cone conjecture holds for every quotient pair $(X/G,\Delta_{G})$ of $(X,\Delta)$ by a finite subgroup $G$ of $\mathrm{Aut}_{k}(X,\Delta)$.
Note that Theorem 1.6 encompasses is a mixed (geometric and arithmetic) equivariant descent result. It generalizes descent results proven in [7] in order to establish analogues of the Kawamata–Morrison cone conjecture for K3 surfaces over arbitrary base fields of characteristics other than 2, and more recently in [50, Section 2.7, Lemma 23] for Enriques surfaces over non-closed fields of characteristic zero. We make note of the following consequence, used notably by [38, Corollary 8.5] with $k$ a function field:
Corollary 1.7.
Let $X$ be a torsor over an abelian variety over a perfect field $k$. Then the movable cone conjecture holds for $X$.
@@
Let $\mathscr{C}$ be a well-clipped cone in a self-dual homogeneous cone $\mathscr{A}$ in $V=V_{\mathbb{Z}}\otimes\mathbb{R}$. Let $\Gamma$ denote the group $\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$. Let $G$ be a subgroup of $\Gamma$ whose action on $V$ has finite orbits. Then, the invariant cone $\mathscr{C}^{G}$ is well-clipped. Furthermore, if there exists a rational polyhedral fundamental domain for the action of $\Gamma$ on $\mathscr{C}^{+}$, then there exists a rational polyhedral fundamental domain for the action of the centralizer $C_{\Gamma}(G)$ on $\mathscr{C}^{G}$$^{+}$.
@@
Acknowledgements. I want to thank A. Höring, H.-Y. Lin, G. Pacienza and L. Wang for discussions in Fall 2022 that convinced me that descending the Morrison–Kawamata cone conjecture under finite quotients was interesting. Although they do not appear in the final paper, examples and arguments from B. Skauli’s Master thesis [44] were quite formative to my understanding of this question. I am also thankful to M. Monti, K. Oguiso, A. Quedo and A. Sarti for their interest in this problem and for their questions on earlier communications, and in particular to G. Pacienza for discussions in Mai 2023 in Nancy, to C. Lehn for his helpful comments on a first draft, to N. Tsakanikas for useful remarks on the symplectic setting during a visit to Lausanne in March 2025, and to S. Filipazzi for sharing an early draft of [38]. I also thank J. Schneider for pointing me to the reference [23] regarding the Néron–Severi representation of absolute Galois groups of perfect fields. I am partially supported by the projects GAG ANR-24-CE40-3526-01 and the DFG funded TRR 195 - Projektnumber@ 286237555.
@@ 2. Preliminaries
If $G$ is a subgroup of $\mathrm{GL}(V)$, we denote by
@@
the invariant subspace. For a cone $\mathscr{C}$ in $V$ that is preserved by $G$, we denote by $\mathscr{C}^{G}:=\mathscr{C}\cap V^{G}$ the invariant cone.
@@ 2.A. Self-dual homogeneous cones
Let $n\geq 3$. Let $q$ be the quadratic form of signature (1,n1)(1,n-1) on $\mathbb{R}^{n}$, and pick $h\in\mathbb{R}^{n}$ such that $q(h)>0$. The non-degenerate open cone
@@
A classification of $\mathbb{R}$-indecomposable self-dual homogeneous cones was achieved by the 1934 result of Jordan–von Neumann–Wigner [18] on the classification of formally real Jordan algebras, and by the 1970ies Koecher–Vinberg theorem stating an equivalence of categories between formally real Jordan algebras and self-dual homogeneous cones (see e.g. [21], [11, Theorem III.2.1 and Section III.3], and Theorem 2.12 for a precise statement). Here we present the classification.
@@ A Jordan algebra $(V,e,\circ)$ is called formally real if for all $x_{1},\ldots,x_{k}\in V$, we have
The next result is due to Koecher and Vinberg, see for instance [21, Theorem 15].
@@ 2.D. Useful facts from birational geometry
This subsection gathers a few facts from birational geometry, which we will mostly use in the proof of Theorem 1.4. Our first lemma is very close to a result of Xu for the cone $\mathrm{Mov}^{\rm e}$ [50, Theorem 24, Remark on Page 26] and Gachet-Lin-Stenger-Wang for the cone $\mathrm{Nef}^{\rm e}$ [12, Proposition 1.6]. We state it and prove it herefor the cones $\mathrm{Mov}^{+}$ and $\mathrm{Nef}^{+}$, under somewhat weaker assumptions than in [50, Theorem 24].
@@
Note that $F:=f^{*}\mathrm{Mov}^{+}(Y)$ is a face of the cone $\mathrm{Mov}^{+}(X)$. By [12, Proposition 3.6], since $(X,\Delta)$ satisfies the movable cone conjecture, the subgroup $\mathrm{Stab}(F)$ of $\mathrm{PsAut}^{*}(X,\Delta)$ that stabilizes the face $F$ acts on $F$ with a rational polyhedral fundamental domain. To conclude, it suffices to check that $\mathrm{Stab}(F)$ is a subgroup of $f^{-1}\circ\mathrm{PsAut}^{*}(Y,\Delta_{Y})\circ f$.
For $g\in\mathrm{Stab}(F)$, note that there exists $\alpha:X\dashrightarrow X^{\prime}$ a small $\mathbb{Q}$-factorial modification such that the relative interiors of $g^{*}f^{*}\mathrm{Nef}(Y)$ and $f^{*}\alpha^{*}\mathrm{Nef}(Y)$ intersect by [4, Theorem 1.2] (see also [12, Proposition 4.10]). By [12, Lemma 4.2], this means that $f\circ g=\alpha\circ f$, thus $g\in f^{-1}\circ\mathrm{PsAut}^{*}(Y,\Delta_{Y})\circ f$ as wished. ∎
@@
Apply [12, Theorem 1.5 (i), (3) $\Rightarrow$ (4)] to the pair $(X,\Delta)$, and restrict to the Mori chambers in $N^{1}(X)$ associated to small $\mathbb{Q}$-factorial modifications of $Y$ (pre-composed with $f$). ∎
@@
The next lemma is inspired by [10, Lemma 4.6], and is therein also credited to C. Casagrande.
@@
so by [12, Lemma 4.2] (see also [20, Lemma 1.5] in dimension 3), the birational map
@@
Note that applying [10, Lemma 4.6] to the identity from $X$ to itself, any other decomposition of $X$ as a product differs by a mere permutation. Thus, by [16, Exercise III.12.6] and by Lemma 2.17, for a pseudoautomorphism $\alpha$ of $X$, there is a permutation $\sigma$ of the integers between $1$ and $r$ and, for every index $i$, a small $\mathbb{Q}$-factorial modification $\alpha_{i}:X_{i}\dashrightarrow X_{\sigma(i)}$, such that
@@
We conclude with a lemma inspired by [32], [25, Section 7, Page 31], and [12, Proof of Theorem 6.1 (iii) (3a), Page 26-27]. Here, one may think of the $(V_{w})_{w\in W}$ as wall hyperplanes corresponding to flops.
@@
Fix $i$. We now prove the nef cone conjecture for $(X_{i},{\alpha_{i}}_{*}\Delta)$. Denote by $\Sigma_{i}$ the sum of the cones of the form ${\beta_{j}^{-1}}^{*}\Pi\cap\mathrm{Nef}(X_{i})$ for $1\leq j\leq s_{i}$. Note that $\Sigma_{i}$ is a itself rational polyhedral: Indeed, by assumption on the wall hyperplanes $(V_{w})$, any intersection of the form $\alpha_{k}^{*}\mathrm{Nef}(X_{k})\cap\Pi$ for $1\leq k\leq r$ is a rational polyhedral cone. If the translates of $\Sigma_{i}$ by $\mathrm{Aut}(X_{i},{\alpha_{i}}_{*}\Delta)$ cover the cone $\mathrm{Amp}(X_{i})$, then [27, Proposition 4.1] concludes. Let $D\in\mathrm{Amp}(X_{i})$. There exists $\beta\in\mathrm{PsAut}(X,\Delta)$ such that $\beta^{*}\alpha_{i}^{*}D\in\Pi$, in particular it belongs to a chamber $\beta_{j}^{*}\alpha_{i}^{*}\mathrm{Amp}(X_{i})$ for some $j$. By [12, Lemma 4.2], this forces
@@ 3.A. Definition and examples of well-clipped cones
In general, Assumption (ii) could be rephrased, in a lattice-theoretic context, by saying that the $(e_{i})_{i\in I}$ are roots of the lattice $V_{\mathbb{Z}}$, see [5, Section 0.1]. However, we prefer to avoid this terminology altogether, as it is often expected of roots to have norm 2-2, whilst our $(e_{i})_{i\in I}$ can have any negative square.
@@
    The nef cone of a smooth del Pezzo surface $\Sigma$ is rational polyhedral, and well clipped. For $\rho(\Sigma)=1$ or $2$, it follows from Item (1). Otherwise, we can take the quadratic form to be the intersection form on $N^{1}(\Sigma)$, and the $(e_{i})_{i\in I}$ in Remark 3.4 to be the finitely many classes of (1)(-1)-curves.
@@
    The movable cone of a smooth projective hyperkähler manifold is well clipped by [30, Lemma 6.22] and [31, Theorem 1.1].
@@
    The movable cone of a projective primitive symplectic variety with terminal $\mathbb{Q}$-factorial singularities is well clipped [25, Theorem 3.10, Lemma 4.6]; [19, Definition 3.1, Theorem 3.11].
@@
    The movable cone of an abelian variety is self-dual homogeneous, thus well clipped [41, Theorem 4.3 and the two paragraphs thereafter].
@@
    The nef cone of $\mathbb{P}^{2}$ blown up at 99 general points and the movable cone of $\mathbb{P}^{3}$ blown up at 8 very general points are both well clipped, by [35] and [45] respectively.
@@
    with $h^{1}(\mathscr{O}_{X_{i}})=0$ for $2\leq i\leq k$ is well clipped if and only if the movable cone of each factor is well clipped. This follows from [16, Exercise III.12.6], Lemmas 2.16 and 3.5.
@@
By contradiction, assume that $E^{2}\leq-3$. Then $K_{S}\cdot E>0$. Since $E$ spans an extremal ray of the Mori cone of $S$, the cone theorem gives an elementary Mori contraction $\varepsilon:S\to S^{\prime}$ with exceptional divisor $E$. Computing discrepancies, one checks that $E$ is contracted to a terminal, hence smooth point of $S^{\prime}$. Castelnuovo’s contraction theorem applies and $E$ has square 1-1 in $S$, contradiction. ∎
@@
    By the Beauville–Bogomolov decomposition [3], either $\chi(\mathscr{O}_{\Sigma})\geq 0$ or $S=\Sigma$ is a bielliptic surface. Lemma 3.7 and the fact that bielliptic surfaces have Picard rank 22 respectively conclude.